The perpendicular height of a triangle is 3 cm less than the
length of its base. The area of the triangle is 27 cm2. Form a
quadratic equation to represent the information and solve to
find the dimensions of the triangle.

Answer by ewatrrr(24785)

A = (1/2)bh
Question States***
(1/2)b(b-3) = 27cm^2
b(b-3) = 54
b^2 - 3b -54 = 0
(b+6)(b-9) = 0 (Tossing out the negative solution for unit measure)
b = 9cm and h = 6cm 9-3
And...checking ***
54/2 = 27